# Real Noncommutative Convexity II: Extremality and nc convex functions

### v2

@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
Given a real nc convex set $K$, we define real maximal, pure and extreme points analogously to their complex counterpart.
@@
A point $x\in K_{n}$ is called pure if when it is written as a finite nc convex combination, $x=\sum_{i}\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ for $x_{i}\in K_{n_{i}}$ and non-zero $\alpha_{i}\in M_{n_{i},n}(\mathbb{R})$ such that $\sum_{i}\alpha_{i}^{*}\alpha_{i}=1_{n}$, then $\alpha_{i}=c_{i}\beta_{i}$ where $\beta_{i}$ is an isometry and $c_{i}$ is a positive number such that $\beta_{i}^{\mathsf{T}}x_{i}\beta_{i}=x$.
Viewing $K=\mathrm{ncS}(V)$ for an appropriate real operator system $V$ as in [8, Section 4], this corresponds to $\psi\leq\varphi$ for a cp map $\psi$, implying that $\psi$ is a positive scalar multiple of $\varphi$. Indeed it is obvious that such a $\varphi$ is pure in the first sense above (note $x=t_{i}\alpha_{i}^{*}x_{i}\alpha_{i}$ evaluated at 1 implies $1=t_{i}\alpha_{i}^{*}\alpha_{i}$, so $\alpha_{i}$ is a multiple of an isometry). Conversely, if $\varphi=\varphi_{1}+\varphi_{2}$ for cp maps $\varphi_{i}$, use [9, Lemma 6.3] to write $\varphi=b_{1}\psi_{1}b_{1}+b_{2}\psi_{2}b_{2}$ for ucp maps $\psi_{i}$ and matrices $b_{i}\geq 0$ with $b_{1}^{2}+b_{2}^{2}=1$. If $\varphi$ is pure in the first sense then $\varphi_{i}=b_{i}\psi_{i}b_{i}$ is a positive multiple of $\varphi$. For most practical purposes this allows us in most situations to drop the requirement that $\alpha_{i}$ is a positive scalar multiple of an isometry above, instead saying that $\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ is a positive scalar multiple of $x$. See Remark 6.1.6 in [14].
@@
A point $x\in K_{n}$ is called extreme if when it is written as a finite nc convex combination, $x=\sum_{i}\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ for $x_{i}\in K_{n_{i}}$ and non-zero $\alpha_{i}\in M_{n_{i},n}(\mathbb{R})$ such that $\sum_{i}\alpha_{i}^{*}\alpha_{i}=1_{n}$, then $\alpha_{i}=c_{i}\beta_{i}$ where $\beta_{i}$ is an isometry and $c_{i}$ is a positive number such that $\beta_{i}^{\mathsf{T}}x_{i}\beta_{i}=x$ and $x_{i}$ commutes with $\beta_{i}\beta_{i}^{\mathsf{T}}$ for each $i$ (so that $x_{i}$ decomposes with respect to the range of $\alpha_{i}$ as $x_{i}=y_{i}\oplus z_{i}$ with $y_{i}$ unitarily equivalent to $x$). The set of real extreme points is denoted by $\partial K$ or Ch(K)(K).
We sometimes write the collection of extreme points as Ch(K)(K), since it corresponds to the Choquet boundary, not the Shilov boundary. The later Corollary 2.14 will say that these are essentially the same as the Arveson boundary representations.
@@
This is a well known result due to Størmer for complex $C^{*}$-algebras, and the real case follows immediately. If $\pi=\frac{1}{2}(\varphi+\psi)$ where $\varphi,\psi\in{\rm UCP}(A,B(H))$ then $\pi_{c}=\frac{1}{2}(\varphi_{c}+\psi_{c})$. Thus $\varphi_{c}=\pi_{c}$ and $\varphi=\pi$. (See also [33, Corollary 3.10] for a complementary result.) ∎
We next prove a useful characterization of the maximal elements of a complexification. We recall the fact from the first three lines of the last paragraph of the proof of [8, Theorem 3.4] that an element $x\in K_{2n}$ of a real nc convex set $K$ is of the form $c(a,b)$ for $a,b\in K_{n}$ if and only if for the unitary matrix $W=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$ we have $W^{*}xW=x$. In this case, $c(u^{*}xu)=x$. Real nc affine maps preserve such elements. Indeed if $\omega$ is nc affine then $W^{*}\omega(c(x))W=\omega(W^{*}c(x)W)=\omega(c(x))$. Hence also $c(u^{*}\omega(c(x))u)=\omega(c(x))$. We use this in the proof below.
@@
Using the the facts before the theorem and that $\psi(v)=c(\varphi(v))$ we see that this is also a ucp extension of $\psi$. Then $\tilde{\rho}_{c}:C^{*}(V)_{c}\to M_{2n}(\mathbb{C})$ is ucp. Thus $u^{*}\tilde{\rho}_{c}(\cdot)u:C^{*}(V)_{c}\to M_{n}(\mathbb{C})$ is ucp, and extends $\varphi$. Indeed writing $c:M_{n}(\mathbb{C})\to M_{2n}(\mathbb{R})$ as $c_{n}$, we have
@@
A direct proof of this result is possible (see e.g. the proof of Lemma 3.3). Instead we deduce it from swiftly from Theorem 2.6. By that result $x+i0\in K_{c}$ is complex maximal if and only if $x\oplus x$ is real maximal in $K$. However it is a pleasant exercise that this is equivalent to $x$ being maximal. ∎
Thus maximal elements or maximal ucp maps behave perfectly in regard to complexification.
One may obtain the following Dritschel-McCullough theorem in the real case by essentially the same proof as in the complex case, e.g. as it is presented in Arveson’s account [3] (an ordinal argument involving direct limits of Hilbert spaces, which includes a contribution by N. Ozawa). This is sketched in an Appendix by Travis Russell. Alternatively, and for its independent interest, we will deduce it instantaneously from the complex case by complexification, as follows:
@@
This gives a (new) route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). Namely the real case of the Dritschel-McCullough-Arveson route (see [18, 3] or as mentioned before Theorem 2.4 or after Theorem 2.17), but using the ‘real arguments’ above.
As in [14] we define the barycenter of a ucp map $\mu:C(K)\to M_{n}$ to be $x$ if $\mu_{|A(K)}=x$. We say $x$ has a unique representing map if $\delta_{x}$ is the unique such ucp $\mu$. Since $C(K)=C^{*}_{\max}(A(K))$, (the real case of) Propositions 5.2.3 and 5.2.4 in [14] are also a rephrasing in the language above of parts of Theorem 2.4, and would need little thought for those familiar with the basics here (although the proofs in [14] work in the real case). Thus:
@@
We say that $x\in K_{n}$ is irreducible if its (selfadjoint) image in $M_{n}$ is irreducible in the sense that every element in $(M_{n})_{\rm sa}$ commuting with $x$ is a scalar multiple of $I_{n}$. This is equivalent to $\delta_{x}$ being an irreducible *-representation. Every pure $x\in K$ is irreducible, as in the complex case [2]. If $K$ is a real nc convex set we say that $x\in K_{n}$ is complex-irreducible if its (selfadjoint) image is irreducible in $M_{n}(\mathbb{C})$.
@@
For $K$ a real nc convex set, $x\in K$ is extreme if and only if it is pure and maximal.
@@
Let $\varphi$ be a ucp map on a real $C^{*}$-algebra $A$.
@@
    $\varphi$ is maximal (resp. extreme) if and only if it is a *-representation (resp.  irreducible *-representation) of $A$.
@@
(1) This is essentially [2, Corollary 1.4.3]. The proof in the real case requires a couple of additional observations. E.g. we need to add to Arveson’s proof the fact about real irreps from [29, Proposition 5.3.7 (1)]. Indeed the proof of [29, Proposition 5.3.7 (1) and Proposition 4.3.4 (3) and (4)] shows that $\varphi(V)^{\prime}$ is a real von Neumann algebra whose projections norm densely span its selfadjoint part. So $\varphi(V)$ is irreducible if and only there are no nontrivial such projections, or equivalently if $(\varphi(V)^{\prime})_{\rm sa}=\mathbb{R}\,I$. We also note that in Arveson’s proof(s) in [2, Section 1.4] we may assume that the operators $T_{i}$ there are selfadjoint, and where he says that $\pi(B)^{\prime}$ is scalar, we use $(\pi(B)^{\prime})_{\rm sa}$ is scalar. For ucp maps Arveson’s usage of ‘pure’ in [2] is equivalent to Definition 2.2.
@@
The proof of the following extension of Corollary 2.9 is the same in the real case (see [14, Theorem 6.1.9]).
@@
For $K$ compact real nc convex, a point $x\in K_{n}$ is extreme if and only if the representation $\delta_{x}:C(K)\to M_{n}$ is both irreducible and is the unique representing map of $x$.
@@
For a real operator system $V$ with $K=\mathrm{ncS}(V)$, the extreme points of $K$ are precisely the restrictions to $V$ of boundary representations of $V$.
@@
Let $V$ be a real operator system and $K$ its associated real nc convex set. Every real pure ucp map $\varphi:V\to B(H)$ can be dilated to an extreme ucp map $\widetilde{\varphi}:\mathcal{S}\to B(L)$. Every real pure in $K$ has an extreme dilation.
@@
Let $V$ be a real operator system and $K$ its associated real nc convex set. If $x\in M_{n}(V)$ for $n<\infty$ then there is an $m$ and an extreme ucp map $\varphi:V\to M_{m}(\mathbb{R})$ with $\|\varphi_{n}(x)\|=\|x\|$. For each $\epsilon>0$ there is also a pure matrix state $\varphi:V\to M_{k}(\mathbb{R})$ for $k<\infty$ with $\|\varphi_{n}(x)\|>\|x\|-\epsilon$.
@@
This is a combination of the real cases of Theorems 3.1, 3.3 and 3.4 in [13]. All of this (including the alternative arguments given there) works the same in the real case, with as usual $\mathbb{C}$ replaced by $\mathbb{R}$, and recalling that all completely positive maps are selfadjoint. One argument there uses [23, Theorem 2.2], which states that for a real operator system $V$, finite $k$, and pure cp map $\psi:V\to M_{k}(\mathbb{R})$ there exists $m\leq k$, pure $\varphi\in{\rm UCP}(V,M_{m}(\mathbb{R}))$ and $\gamma\in M_{m,k}(\mathbb{R})$ with $\psi=\gamma^{\mathsf{T}}\varphi(\cdot)\gamma$. In turn this uses in the second paragraph of its proof some complex finite dimensional linear algebra, and some tricks that often fail for real spaces. Because of its independent interest we check this in the real case. Let $b=\psi(1)$ and $p$ the projection onto Ran $b$. To prove in that second paragraph that $p\in\psi(V)^{\prime}$ we proceed as follows. Consider completely positive $\psi_{c}:V_{c}\to M_{k}(\mathbb{C})$, and view $p,b\in M_{k}(\mathbb{R})\subset M_{k}(\mathbb{C})$. Indeed $p$ is then the projection in $M_{k}(\mathbb{C})$ onto $\mathbb{C}^{k}$. The proof in that second paragraph shows that $p\in\psi_{c}(V_{c})^{\prime}$, so that $p\in\psi(V)^{\prime}$. ∎
We also of course now immediately obtain [14, Theorem 6.2.4] in the real case.
@@
For $K$ a compact real nc convex set, define the Hilbert space $H=\oplus_{n}\oplus_{x\in(\partial K)_{n}}\,H_{n}$ and $\pi:C(K)\to B(H)$ the representation $\pi=\oplus_{x\in\partial K}\delta_{x}$. Then, the restriction $\pi|_{A(K)}$ is a ucoi and $\pi(C(K))$ is *-isomorphic to $C^{*}_{\rm min}(A(K))$.
Just as in the complex case, the $C^{*}$-algebra generated by the image of $V$ under a direct sum of sufficiently many extreme ucp maps (or boundary representations), is (a copy of) the $C^{*}$-envelope of $V$. Indeed this direct sum is maximal [14, Corollary 3.5]. Thus, a quick method to construct the real $C^{*}$-envelope of a real operator system $V$ is thus to take a maximal dilation $\rho$ of an inclusion $V\subset B(H)$. The $C^{*}$-algebra generated by $\rho(V)$ is the real $C^{*}$-envelope. (Note also that $\rho$ is maximal in $K_{c}$ by Corollary 2.7, so the $C^{*}$-algebra generated by $\rho_{c}(V_{c})=\rho(V)+i\rho(V)$ is a copy of the complex $C^{*}$-envelope of $V_{c}$.)
@@
The proof is the same as in the complex case, following quickly from Theorem 2.17 and [8, Theorem 3.6]. See also the real matrix Krein-Milman theorem in [22]. The following is the real case of Davidson and Kennedy’s noncommutative analogue of Milman’s partial converse to the Krein-Milman theorem [14, Theorem 6.4.3]:
@@
For the real case, one needs to know that various properties from [17] of irreducible representations and corresponding GNS representations used in [14, Theorem 6.4.3], as well as Proposition 3.4.2 (ii) in [17], all hold in the real case. The former properties are checked partially in [29] (see also [9, Section 5]). The latter result appeals to [17, Lemma 3.4.1], which in turn relies on facts on extreme functionals from Proposition 2.5.5 there. The latter is well known in the real case (see [29, Proposition 5.2.7] and its proof). Otherwise the proof proceeds almost exactly as in [14, Theorem 6.4.3]. For the readers convenience additional details are given in the Appendix below by T. Russell. ∎

### v3

@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
Given a real nc convex set $K$, we define real maximal, pure and nc extreme points analogously to their complex counterpart.
@@
A point $x\in K_{n}$ is called pure if when it is written as a finite nc convex combination, $x=\sum_{i}\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ for $x_{i}\in K_{n_{i}}$ and non-zero $\alpha_{i}\in M_{n_{i},n}(\mathbb{R})$ such that $\sum_{i}\alpha_{i}^{\mathsf{T}}\alpha_{i}=1_{n}$, then $\alpha_{i}=c_{i}\beta_{i}$ where $\beta_{i}$ is an isometry and $c_{i}$ is a positive number such that $\beta_{i}^{\mathsf{T}}x_{i}\beta_{i}=x$.
If $K=\mathrm{ncS}(V)$ for an appropriate real operator system $V$ as in [8, Section 4], this corresponds to $\psi\leq\varphi$ for a cp map $\psi$, implying that $\psi$ is a positive scalar multiple of $\varphi$. Indeed it is obvious that such a $\varphi$ is pure in the first sense above (note $x=t_{i}\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ evaluated at 1 implies $1=t_{i}\alpha_{i}^{\mathsf{T}}\alpha_{i}$, so $\alpha_{i}$ is a multiple of an isometry). Conversely, if $\varphi=\varphi_{1}+\varphi_{2}$ for cp maps $\varphi_{i}$, use [9, Lemma 6.3] to write $\varphi=b_{1}\psi_{1}b_{1}+b_{2}\psi_{2}b_{2}$ for ucp maps $\psi_{i}$ and matrices $b_{i}\geq 0$ with $b_{1}^{2}+b_{2}^{2}=1$. If $\varphi$ is pure in the first sense then $\varphi_{i}=b_{i}\psi_{i}b_{i}$ is a positive multiple of $\varphi$. This allows us in many situations (such as if the embedding of $K$ is ‘nc regular’, or simply if $K$ lies in a nc hyperplane not containing 0, as in Remark 6.1.6 in [16]) to drop the requirement that $\alpha_{i}$ is a positive scalar multiple of an isometry above, instead saying that $\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ is a positive scalar multiple of $x$.
@@
A point $x\in K_{n}$ is called (nc) extreme if when it is written as a finite nc convex combination, $x=\sum_{i}\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ for $x_{i}\in K_{n_{i}}$ and non-zero $\alpha_{i}\in M_{n_{i},n}(\mathbb{R})$ such that $\sum_{i}\alpha_{i}^{\mathsf{T}}\alpha_{i}=1_{n}$, then $\alpha_{i}=c_{i}\beta_{i}$ where $\beta_{i}$ is an isometry and $c_{i}$ is a positive number such that $\beta_{i}^{\mathsf{T}}x_{i}\beta_{i}=x$ and $x_{i}$ commutes with $\beta_{i}\beta_{i}^{\mathsf{T}}$ for each $i$ (so that $x_{i}$ decomposes with respect to the range of $\alpha_{i}$ as $x_{i}=y_{i}\oplus z_{i}$ with $y_{i}$ unitarily equivalent to $x$). The set of real extreme points is denoted by $\partial K$ or Ch(K)(K).
We sometimes write the collection of nc extreme points as Ch(K)(K), since it corresponds to the Choquet boundary, not the Shilov boundary. The later Corollary 2.14 will say that these are essentially the same as the Arveson boundary representations.
@@
This is a well known result due to Størmer for complex $C^{*}$-algebras, and the real case follows immediately. If $\pi=\frac{1}{2}(\varphi+\psi)$ where $\varphi,\psi\in{\rm UCP}(A,B(H))$ then $\pi_{c}=\frac{1}{2}(\varphi_{c}+\psi_{c})$. Thus $\varphi_{c}=\pi_{c}$ and $\varphi=\pi$. (See also [36, Corollary 3.10] for a complementary result.) ∎
We next prove a useful characterization of the maximal elements of a complexification. We recall the fact from the first three lines of the last paragraph of the proof of [8, Theorem 3.4] that an element $x\in K_{2n}$ of a real nc convex set $K$ is of the form $c(a,b)$ for $a,b\in K_{n}$ if and only if for the unitary matrix $W=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$ we have $W^{*}xW=x$. In this case, $c(u^{*}xu)=x$. Real nc affine maps preserve such elements. Indeed if $\omega$ is nc affine then $W^{*}\omega(x)W=\omega(W^{*}xW)=\omega(x)$ if $x=c(a,b)$. Hence also $c(u^{*}\omega(x)u)=\omega(x)$. We use this in the proof below.
@@
Using the facts before the theorem and that $\psi(v)=c(\varphi(v))$ we see that this is also a ucp extension of $\psi$. Then $\tilde{\rho}_{c}:C^{*}(V)_{c}\to M_{2n}(\mathbb{C})$ is ucp. Thus $u^{*}\tilde{\rho}_{c}(\cdot)u:C^{*}(V)_{c}\to M_{n}(\mathbb{C})$ is ucp, and extends $\varphi$. Indeed writing $c:M_{n}(\mathbb{C})\to M_{2n}(\mathbb{R})$ as $c_{n}$, we have
@@
A direct proof of this result is possible (e.g. this is similar to the proof of a special case of Theorem 2.25 in an earlier ArXiV draft). Instead we deduce it swiftly from Theorem 2.6. By that result $x+i0\in K_{c}$ is complex maximal if and only if $x\oplus x$ is real maximal in $K$. However it is a pleasant exercise that this is equivalent to $x$ being maximal. ∎
Thus maximal elements or maximal ucp maps behave perfectly in regard to complexification. (Confirming this, we shall show at the end of the section that if $K$ a complex compact nc convex set and $x\in K$, then $x$ is real maximal in $K$ if and only if $x$ is complex maximal in $K$. This is not true for extremal elements.)
One may obtain the following Dritschel-McCullough theorem in the real case by essentially the same proof as in the complex case, e.g. as it is presented in Arveson’s account [3] (an ordinal argument involving direct limits of Hilbert spaces, which includes a contribution by N. Ozawa). This was sketched in a previous version of our paper, in the Appendix by Travis Russell. We removed this since it is essentially identical to Arveson’s account, and since the latter is now published in [14]. Instead, or alternatively, and for its independent interest, we will deduce it instantaneously from the complex case by complexification, as follows:
@@
This gives a (new) route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). Namely the real case of the Dritschel-McCullough-Arveson route (see [20, 3] or [14, Theorem 14.9.12], or as mentioned before Theorem 2.4 or after Theorem 2.17), but using the ‘real arguments’ above. For example, if we take a maximal dilation $\rho$ of the inclusion $V\subset B(H)$, using our swift Theorem 2.8, then the $C^{*}$-algebra generated by $\rho(V)$ is the real $C^{*}$-envelope.
As in [16] we define the barycenter of a ucp map $\mu:C(K)\to M_{n}$ to be $x\in K$ if $\mu_{|A(K)}$ equals evaluation at $x$. We say $x$ has a unique representing map if $\delta_{x}$ is the unique such ucp $\mu$. Since $C(K)=C^{*}_{\max}(A(K))$, (the real case of) Propositions 5.2.3 and 5.2.4 in [16] are also a rephrasing in the language above of parts of Theorem 2.4, and would need little thought for those familiar with the basics here (and the proofs in [16] work in the real case). Thus:
@@
We say that $x\in K_{n}$ is irreducible if its (selfadjoint) image in $M_{n}$ (that is, $\{f(x):f\in A(K)\}$) is irreducible in the sense that every element in $(M_{n})_{\rm sa}$ commuting with $x$ is a scalar multiple of $I_{n}$. This is equivalent to $\delta_{x}$ being an irreducible *-representation. Every pure $x\in K$ is irreducible, as in the complex case [2]. If $K$ is a real nc convex set we say that $x\in K_{n}$ is complex-irreducible if its (selfadjoint) image $\{f(x):f\in A_{\mathbb{R}}(K)\}$ is irreducible as a subset of $M_{n}(\mathbb{C})$.
As shown in Examples such as 3.6 below, a point $x\in K$ for real nc convex $K$ may be pure but $x\in K_{c}$ may not be. Similarly, by the statement after Lemma 3.7, a point $x\in K_{c}$ may be extreme while $c(x)$ is not. Thus we cannot prove the Davidson-Kennedy theorem (see Theorem 6.2.2 in [16] or 14.10.8 in [14]) via complexification, nor can we use complexification to deduce the existence of extreme points. Therefore, because pure (i.e. irreducible) points do not complexify well, we cannot do a proof of the real nc Krein-Milman theorem analogous to how we did Theorem 2.8. Thus, we now spend time to prove the real nc Krein-Milman theorem directly.
@@
For $K$ a real nc convex set, $x\in K$ is nc extreme if and only if it is pure and maximal.
@@
Let $\varphi:A\to M_{n}$ be a ucp map on a real $C^{*}$-algebra $A$.
@@
    $\varphi$ is maximal (resp. nc extreme) if and only if it is a *-representation (resp. irreducible *-representation) of $A$.
@@
(1) This is essentially [2, Corollary 1.4.3]. The proof in the real case requires a couple of additional observations. E.g. we need to add to Arveson’s proof the fact about real irreps from [32, Proposition 5.3.7 (1)]. Indeed the proof of [32, Proposition 5.3.7 (1) and Proposition 4.3.4 (3) and (4)] shows that $\varphi(V)^{\prime}$ is a real von Neumann algebra whose projections norm densely span its selfadjoint part. So $\varphi(V)$ is irreducible if and only there are no nontrivial such projections, or equivalently if $(\varphi(V)^{\prime})_{\rm sa}=\mathbb{R}\,I$. We also note that in Arveson’s proof(s) in [2, Section 1.4] we may assume that the operators $T_{i}$ there are selfadjoint, and where he says that $\pi(B)^{\prime}$ is scalar, we use $(\pi(B)^{\prime})_{\rm sa}$ is scalar. For ucp maps Arveson’s usage of ‘pure’ in [2] is equivalent to Definition 2.2.
@@
The proof of the following extension of Corollary 2.9 is the same in the real case (see [16, Theorem 6.1.9]).
@@
For $K$ compact real nc convex, a point $x\in K_{n}$ is nc extreme if and only if the representation $\delta_{x}:C(K)\to M_{n}$ is both irreducible and is the unique representing map of $x$.
@@
For a real operator system $V$ with $K=\mathrm{ncS}(V)$, the nc extreme points of $K$ are precisely the restrictions to $V$ of boundary representations of $V$.
@@
Let $V$ be a real operator system and $K$ its associated real nc convex set. Every real pure ucp map $\varphi:V\to B(H)$ can be dilated to a nc extreme ucp map $\widetilde{\varphi}:\mathcal{S}\to B(L)$. Every real pure in $K$ has an extreme dilation.
@@
Let $V$ be a real operator system and $K$ its associated real nc convex set. If $x\in M_{n}(V)$ for $n<\infty$ then there is an $m$ and a nc extreme ucp map $\varphi:V\to M_{m}(\mathbb{R})$ with $\|\varphi_{n}(x)\|=\|x\|$. For each $\epsilon>0$ there is also a pure matrix state $\varphi:V\to M_{k}(\mathbb{R})$ for $k<\infty$ with $\|\varphi_{n}(x)\|>\|x\|-\epsilon$.
@@
This is a combination of the real cases of Theorems 3.1, 3.3 and 3.4 in [15]. All of this (including the alternative arguments given there) works the same in the real case, with as usual $\mathbb{C}$ replaced by $\mathbb{R}$, and recalling that all completely positive maps are selfadjoint. One argument there uses [25, Theorem 2.2], which states that for a real operator system $V$, finite $k$, and pure cp map $\psi:V\to M_{k}(\mathbb{R})$ there exists $m\leq k$, pure $\varphi\in{\rm UCP}(V,M_{m}(\mathbb{R}))$ and $\gamma\in M_{m,k}(\mathbb{R})$ with $\psi=\gamma^{\mathsf{T}}\varphi(\cdot)\gamma$. In turn this uses in the second paragraph of its proof some complex finite dimensional linear algebra, and some tricks that often fail for real spaces. Because of its independent interest we check this in the real case. Let $b=\psi(1)$ and $p$ the projection onto Ran $b$. To prove in that second paragraph that $p\in\psi(V)^{\prime}$ we proceed as follows. Consider completely positive $\psi_{c}:V_{c}\to M_{k}(\mathbb{C})$, and view $p,b\in M_{k}(\mathbb{R})\subset M_{k}(\mathbb{C})$. Indeed $p$ is then the projection in $M_{k}(\mathbb{C})$ onto $\mathbb{C}^{k}$. The proof in that second paragraph shows that $p\in\psi_{c}(V_{c})^{\prime}$, so that $p\in\psi(V)^{\prime}$. ∎
We also of course now immediately obtain [16, Theorem 6.2.4] in the real case.
@@
For $K$ a compact real nc convex set, define the Hilbert space $H=\oplus_{n}\oplus_{x\in(\partial K)_{n}}\,H_{n}$ and $\pi:C(K)\to B(H)$ the representation $\pi=\oplus_{x\in\partial K}\delta_{x}$. Then, the restriction $\pi|_{A(K)}$ is a unital complete order isomorphism and $\pi(C(K))$ is *-isomorphic to $C^{*}_{\rm min}(A(K))$.
Just as in the complex case, the $C^{*}$-algebra generated by the image of $V$ under a direct sum of sufficiently many extreme ucp maps (or boundary representations), is (a copy of) the $C^{*}$-envelope of $V$. Indeed this direct sum is maximal [16, Corollary 3.5]. Thus, a quick method to construct the real $C^{*}$-envelope of a real operator system $V$ in $B(H)$ is to take a maximal dilation $\rho$ of the inclusion $V\subset B(H)$. The $C^{*}$-algebra generated by $\rho(V)$ is the real $C^{*}$-envelope. (Note also that $\rho$ is maximal in $K_{c}$ by Corollary 2.7, so the $C^{*}$-algebra generated by $\rho_{c}(V_{c})=\rho(V)+i\rho(V)$ is a copy of the complex $C^{*}$-envelope of $V_{c}$.)
@@
The proof is the same as in the complex case, following quickly from Theorem 2.17 and [8, Theorem 3.6]. See also the real matrix Krein-Milman theorem in [24]. The following is the real case of Davidson and Kennedy’s noncommutative analogue of Milman’s partial converse to the Krein-Milman theorem [16, Theorem 6.4.3]:
@@
For the real case, one needs to know that various properties from [19] of irreducible representations and corresponding GNS representations used in [16, Theorem 6.4.3], as well as Proposition 3.4.2 (ii) in [19], all hold in the real case. The former properties are checked partially in [32] (see also [9, Section 5]). The latter result appeals to [19, Lemma 3.4.1], which in turn relies on facts on extreme functionals from Proposition 2.5.5 there. The latter is well known in the real case (see [32, Proposition 5.2.7] and its proof). Otherwise the proof proceeds almost exactly as in [16, Theorem 6.4.3]. For the readers convenience additional details are given in the Appendix below by T. Russell. ∎